OpenStudy (anonymous):

Write the direct variation of the equation if y varies directly with x and y = 45 when x = 9.

OpenStudy (luigi0210):

Direct variation equation: y=kx

OpenStudy (luigi0210):

So it's pretty much saying to plug in values, and solve for k

OpenStudy (anonymous):

Since, y varies directly with x Mathematically it can be written as \[y \alpha x \rightarrow y=kx...(k=constant)\] Now at y = 45 and x = 9, k=? \[45 = k \times 9 \rightarrow k= \frac{45}{9} \rightarrow k=5\] substituting k=5 in y=kx we find y=5x, which is the required equation of the direct variation.

OpenStudy (anonymous):

@francecsamusco